A random variable X has the cumulative distribution function given as

 x 2 − 2 x + 2
 1, for x ≥ 2
0, F(x)= 2
for x < 1
, for1≤x<2
Calculate the variance of X.
1
Expert's answer
2020-05-18T18:46:52-0400
"F(x) = \begin{cases}
 0 &\text{for } x<1 \\
 \dfrac{x^2-2x+2}{2} &\text{for } 1\leq x<2 \\
 1 &\text{for } x\geq2
\end{cases}"

The magnitude of the jump in the CDF that occur at "x=1" is "\dfrac{1}{2}."

"f(x)=F'(x) = \begin{cases}
 
 {\delta(x-1)\over 2}+x-1 &\text{for } 1\leq x<2 \\
 0 &\text{otherwise } 
\end{cases}"

"E(X)=\displaystyle\int_{1}^2x( {\delta(x-1)\over 2}+x-1)dx="

"={x\over 2}|_{x=1}+\big[{x^3\over 3}-{x^2\over 2}\big]\begin{matrix}
 2 \\
 1
\end{matrix}"

"={1\over 2}+{8\over 3}-2-{1\over 3}+{1\over 2}={4\over 3}"

"E(X^2 )=\displaystyle\int_{1}^2x^2 ( {\delta(x-1)\over 2}+x-1)dx="

"={x^2\over 2}|_{x=1}+\big[{x^4\over 4}-{x^3\over 3}\big]\begin{matrix}
 2 \\
 1
\end{matrix}="


"={1\over 2}+4-{8\over 3}-{1\over 4}+{1\over 3}={23\over 12}"

"V(X)=E(X^2)-(E(X))^2={23\over 12}-({4\over 3})^2={5\over 36}"